### v5

# QCPINN: Quantum-Classical Physics-Informed Neural Networks for Solving PDEs

Abstract
Physics-informed neural networks (PINNs) have emerged as promising methods for solving partial differential equations (PDEs) by embedding physical laws within neural architectures. However, these classical approaches often require a large number of parameters to achieve reasonable accuracy, particularly for complex PDEs. In this paper, we present a quantum-classical physics-informed neural network (QCPINN) that combines quantum and classical components, allowing us to solve PDEs with significantly fewer parameters while maintaining comparable accuracy and convergence to classical PINNs. We systematically evaluated two quantum circuit architectures across various configurations on five benchmark PDEs to identify optimal QCPINN designs. Our results demonstrate that the QCPINN achieves stable convergence and comparable accuracy, while requiring approximately 10% of the trainable parameters used in classical approaches. It also results in a 40% reduction in the relative error $L_{2}$ for the convection-diffusion equation. These findings demonstrate the potential of parameter efficiency as a measurable quantum advantage in physics-informed machine learning, significantly reducing model complexity while preserving solution quality. This approach presents a promising solution to the computational challenges associated with solving PDEs.
1 Introduction
Recent advancements in machine learning have transformed approaches to solving PDEs, with PINNs emerging as an innovative technique. PINNs embed physical laws directly into neural network architectures, eliminating the need for labeled training data while strictly enforcing physical constraints. This approach has demonstrated considerable success in modeling various physical phenomena across scientific and engineering domains [1, 2, 3]. Although they hold great potential, research is underway to improve the accuracy of PINNs in modeling different PDEs and physical systems. They often face challenges with stiff systems, capturing steep gradients, and efficiently navigating high-dimensional parameter spaces. While increasing model expressivity through additional parameters can alleviate these challenges, this approach introduces new problems: increased computational cost, risk of overfitting, and unstable convergence behavior during training [4, 5]. Such a trade-off between model complexity, generalization, and convergence ability is a fundamental challenge in neural network design.
Quantum computing presents a promising paradigm for overcoming these limitations. With its ability to process high-dimensional data and exploit quantum principles such as superposition and entanglement, quantum algorithms offer unique advantages for solving complex computational problems. For instance, the Harrow-Hassidim-Lloyd (HHL) algorithm has demonstrated theoretical exponential speedups for specific linear systems, while variational quantum algorithms provide practical near-term solutions [6, 7, 8]. Moreover, recent developments [9, 10, 11] indicate that quantum neural networks possess an expressive power that could significantly enhance the learning capabilities for modeling complex physical systems.
Although both PINNs and quantum computing have advanced independently, their integration remains largely unexplored. Several preliminary efforts have emerged: Sedykh et al. proposed HQPINN [12], focusing on computational fluid dynamics in 3D Y-shaped mixers; Dehaghani et al. introduced QPINN [13], integrating a dynamic quantum circuit with classical computing methodologies for optimal control problems; Trahan et al. [14] investigated both purely quantum and hybrid PINNs, highlighting parameter efficiency but limiting application to parametrized quantum circuit (PQC) networks with simple PDEs. Leong et al. proposed HQPINN [15], which is inspired by Fourier neural operators. However, existing research has not comprehensively established whether quantum-enhanced models can improve upon purely classical PINNs across a broad range of PDEs efficiently and reliably.
We aim to investigate the limitations of traditional PINNs by integrating quantum computing circuits with classical neural networks. We propose that such an integrated or hybrid approach can leverage complementary strengths: classical computing’s robustness and quantum computing’s enhanced expressivity and inherent parallelism. We believe this method will allow us to solve a wide range of PDEs with similar or better accuracy and convergence rates compared to classical PINNs, while requiring significantly fewer trainable parameters than classical models.
To this end, we introduce a hybrid architecture named Quantum-Classical Physics-Informed Neural Network (QCPINN). To search for a suitable QCPINN architecture and its configurations, we systematically evaluate multiple dimensions of the design space: two quantum circuit architectures: (a) discrete-variable (DV) qubit-based circuits with configurations, such as four circuit topologies: alternate, cascade, cross-mesh, and layered; and two embedding schemes: amplitude and angle; (b) continuous-variable (CV) circuits with configurations, including two measurement schemes: number and quadrature; three nonlinear operations: Kerr, cubic phase, and cross-Kerr; and two parameterizations: phase-free (magnitude) and full (magnitude and phase). We evaluate the performance of these QCPINN models across various PDEs, such as the Helmholtz and lid-driven cavity equations, while also extending our analysis to the Klein-Gordon, wave, and convection-diffusion equations. Our comprehensive assessment compares (a) solution accuracy, (b) convergence rates, and (c) parameter efficiency between classical PINNs and QCPINN variants.
To ensure reproducibility and facilitate broader adoption of our methods, we have made all components of QCPINN open-source code and publicly available on GitHub at https://github.com/afrah/QCPINN. We hope this repository will enable community engagement and collaborative advancement of this emerging field.
2 Background
2.1 Physics Informed Neural Network (PINNs)
PINNs [16] provide a framework to embed governing physical laws directly into neural network architectures. The PINN loss function combines terms representing governing equations and boundary conditions designed to minimize deviations from physical laws. More formally, the PINN loss function can be defined as:
$\begin{split}\mathcal{L}(\theta)=&\arg\min_{\theta}\sum_{k=1}^{n}\lambda_{k}\mathcal{L}_{k}(\theta)\\ =&\arg\min_{\theta}\left(\lambda_{1}\mathcal{L}_{1}(D[u_{\theta}(\mathbf{x});\mathbf{a}]-f(\mathbf{x}))+\sum_{k=2}^{n}\lambda_{k}\mathcal{L}_{k}(B[u_{\theta}(\mathbf{x})]-g_{k}(\mathbf{x}))\right),\end{split}$ (1)
where, $n$ denotes the total number of loss terms, $\lambda_{k}$ are the weighting coefficients, and $\theta$ represents trainable parameters. Here, $D[u_{\theta}(\mathbf{x});\mathbf{a}]$ and $B[u_{\theta}(\mathbf{x})]$ are the arbitrary differential operators and initial/boundary conditions, respectively, that are applied to the output of the neural network $u_{\theta}(\mathbf{x})$. While the initial and boundary conditions ensure that the network output satisfies predefined constraints at specific spatial or temporal locations, the physics loss acts as a regularizer, guiding the network to learn physically consistent solutions and preventing overfitting to sparse or noisy data. The balance between these loss components is crucial for achieving accurate solutions, as improper weighting can lead to underfitting the initial/boundary conditions or failing to capture the system dynamics correctly.
2.2 Continuous Variable (CV) QNN
In photonic quantum computing, CV computation addresses quantum states with continuous degrees of freedom, such as the position ($x$) and momentum ($p$) quadratures of the electromagnetic field. The computational space is, in principle, unbounded, involving infinite-dimensional Hilbert spaces with continuous spectra. To make these simulations computationally feasible, a cutoff dimension is employed to limit the number of basis states used for approximation. The computational state space is defined as $r=n^{m}$, where $m$ represents the number of qumodes and $n$ is the cutoff dimension. For example, with three qumodes and a cutoff dimension of 2, the state space is 8-dimensional. Ideally, higher cutoff dimensions (up to a certain threshold) yield more precise results but at the cost of increased computational complexity.
In this work, we adhere to the affine transformation introduced in [17]. The proposed circuit ansatz consists of a sequence of operations that manipulate quantum states in the phase space. Each layer of the CV-circuit QNN includes the following key components:
  1. 1.
    Interferometers ($U_{1}(\psi_{1},\phi_{1})$ and $U_{2}(\psi_{2},\phi_{2})$): Linear optical interferometers perform orthogonal transformations, represented by symplectic matrices. These transformations adjust the state by applying rotation and mixing between modes, corresponding to beam splitting and phase-shifting operations.
  2. 2.
    Squeezing Gates ($S(r,\phi_{s})$): These gates scale the position and momentum quadratures independently, introducing anisotropic adjustments to the phase space. Encoding sharp gradients using squeezing gates is particularly appealing for PDE solvers, as it allows fine-tuning of local characteristics [12].
  3. 3.
    Displacement Gates ($D(\alpha,\phi_{d})$): These gates shift the state in phase space by adding a displacement vector $\alpha$, translating the state globally without altering its internal symplectic structure.
  4. 4.
    Non-Gaussian Gate ($\Phi(\lambda)$): To introduce nonlinearity, a non-Gaussian operation, such as a cubic phase, Kerr, and cross-Kerr gates, is applied. This step enables the neural network to approximate nonlinear functions and enhances the ability to solve complex, nontrivial PDEs.
These components together mimic the classical affine transformation such that:
[…]
where, $M$ represents the combined symplectic matrix of the interferometers and squeezing gates, $x$ is the input vector in the phase space, $b$ is the displacement vector, and $\Phi$ is the nonlinear transformation.
In addition to the building blocks above, CV-circuit QNNs offer flexibility in measurement schemes and parameterization strategies. Measurement can be performed in either the number basis, which projects the state onto discrete photon-number states, or the quadrature basis, which yields continuous-valued outcomes corresponding to position or momentum observables. Furthermore, the parameterization of unitary operations may be either full (including both magnitude and phase) or restricted to magnitude-only (by setting phase parameters $\phi$ to zero), which can improve trainability and mitigate optimization challenges at the cost of reduced expressivity.
Although such a model exhibits intriguing theoretical properties, their practical implementation in real-world problem solving seems to suffer from fundamental stability issues. This is mainly due to their inherent sensitivity and reliance on continuous degrees of freedom, which complicate optimization compared to DV-based approaches (discussed in detail in section 2.3). This sensitivity stems from several factors, including the precision required for state preparation, noise from decoherence, and the accumulation of errors in variational quantum circuits. Furthermore, the transformations involved in CV quantum circuits are prone to numerical instabilities, particularly when optimized using gradient-based methods, as seen in the following section. Although foundational studies about the barren plateau problem, such as [18, 19, 20], primarily address DV circuit systems, the analysis indicates that similar issues can arise in CV quantum systems. In particular, the infinite-dimensional Hilbert space and the continuous nature of parameterization can exacerbate barren-plateau effects.
2.3 Discrete Variable (DV) QNN
DV-circuit quantum models do not inherently provide affine transformations as a core feature like CV-based QNNs. However, they can achieve affine-like behaviour through parameterized single-qubit rotations, controlled operations, and phase shifts for amplitude adjustments. Therefore, the affine transform of equation (2) can be reinterpreted to align with their architecture such that:
  1. 1.
    $\mathbf{Mx+b}$: can be achieved through a combination of parameterized quantum gates:
    - $U_{1},U_{2}$: A unitary operator that can be represented by parameterized single-qubit rotations, such as $R_{x}(\theta),R_{y}(\theta)$, or $R_{z}(\theta)$.
    - $S$: A sequence of controlled gates (e.g., CNOT and CZ) responsible for introducing entanglement, which enables the representation of correlated variables and interactions.
    - $D$: Data encoding circuits that map classical data into quantum states through techniques such as basis embedding, amplitude embedding, or angle embedding.
  2. 2.
    Non-linear Activation ($\Phi$): DV-circuit QNNs lack intrinsic quantum nonlinearity due to the linearity of quantum mechanics. However, classical measurements or a combination of classical measurements and non-linear activation functions (such as $Tanh$) can be applied to the output between quantum layers [21, 22].
3 Related Work
In the domain of PDE solvers, quantum algorithms can be broadly categorized into two primary approaches: (a) exclusively quantum algorithms and (b) hybrid quantum-classical methods.
3.1 Exclusively Quantum Algorithms
Exclusively quantum algorithms for solving PDEs typically leverage unitary operations and quantum state embedding techniques to achieve a computational advantage. In contrast to neural networks, exclusively quantum algorithms use fixed quantum circuit structures with no trainable parameters. The celebrated Harrow-Hassidim-Lloyd (HHL) algorithm is an example of this category, which provides exponential speedups to solve linear systems under specific conditions [23]. Extensions of the HHL algorithm have been applied to solve sparse linear systems derived from discretized linear [24, 25] or nonlinear [26] PDEs, leveraging numerical techniques such as linear multistep method [27], Chebyshev pseudospectral method [28], and Taylor series approximations [29]. In fluid dynamics applications, lattice-gas quantum models have been examined for directly solving PDEs [30, 31] or PDEs derived from ODEs through methods such as the Quantum Amplitude Estimation Algorithm (QAEA) [32, 33].
Despite these advancements, purely quantum algorithms for solving PDEs encounter practical challenges such as requirements for quantum error correction [34, 35, 36], limitations in circuit depth [37, 31, 38], and bottlenecks in data embedding [39, 40, 41]. These obstacles, among others, limit the near-term viability of exclusively quantum PDE solvers on today’s noisy intermediate-scale quantum (NISQ) devices.
3.2 Hybrid Quantum-Classical Approaches
Hybrid quantum-classical approaches have emerged as promising solutions to overcome the limitations of exclusively quantum algorithms. These methods combine the robustness, maturity, and stability of classical computation with the enhanced capacity, expressivity, and inherent parallelism of quantum systems. While classical components can efficiently handle pre-/post-processing, optimization, and error stabilization, quantum layers can capture intricate patterns, solve sub-problems, and accelerate specialized computations.
Within this hybrid paradigm, Quantum Neural Networks (QNNs) have demonstrated universal approximation capabilities for continuous functions when implemented with sufficient depth, width, and well-designed quantum feature maps [42, 43, 44]. For instance, Salinas et al. [45] showed that single-qubit networks with iterative input reuploading can approximate bounded complex functions, while Goto et al. [46] established universal approximation theorems for quantum feedforward networks. Schuld et al. [9] revealed that the outputs of parameterized quantum circuits (with suitable data embedding and circuit depth) can be expressed as truncated Fourier series, thereby providing theoretical support for the universal approximation ability of QNNs. Liu et al. [47] provided formal evidence of quantum advantage in function approximation. These developments not only support applications in solving PDEs and optimization tasks but also extend to other computational tasks such as quantum chemistry [48], material sciences [49, 50], and pattern recognition [51], to name a few.
A well-known example of this integrated approach is the Variational Quantum Algorithms (VQA), also known as Variational Quantum Circuits (VQC). These algorithms employ parameterized quantum circuits with tunable gates optimized through classical feedback loops. VQA implementations for solving PDEs include the Variational Quantum Eigensolver (VQE)[52, 53], Variational Quantum Linear Solver (VQLS)[54], and Quantum Approximate Optimization Algorithm (QAOA)[55], as well as emerging methods like Differentiable Quantum Circuits (DQCs)[56, 57], quantum kernel methods [58], and Quantum Nonlinear Processing Algorithms (QNPA) [59].
The aforementioned methods primarily utilize qubit-based quantum computing, which represents the predominant paradigm in quantum computing literature. However, continuous-variable quantum computing (CVQC) [60] offers an alternative computational model that leverages continuous degrees of freedom, such as quadratures and electromagnetic field momentum as discussed in section 2.2. For instance, in [61], Gaussian and non-Gaussian gates were used to manage continuous degrees of freedom, offering enhanced scalability and expressivity for continuous systems. For consistency, we refer to qubit-based models as discrete-variable (DV) circuit models throughout this paper.
Recently, VQA methods have been extended to Physics-informed quantum neural networks (PIQNN) [62, 13, 14, 12, 15], broadening their applicability across physics and engineering domains. These hybrid methods offer several distinct advantages when implemented within the PINN framework.
Firstly, these models enable the flexible configuration of inputs and outputs independent of qubit/qumode counts or quantum layer depth, making them particularly effective for scenarios where the input dimension (e.g., temporal and spatial coordinates) differs from output dimensions (e.g., dependent variables like velocity, pressure, and force).
Secondly, hybrid approaches facilitate increased trainable parameter counts without the need for deeper quantum circuits, additional qubits, or higher cutoff dimensions in CV quantum systems. These factors are particularly crucial in the NISQ era, where hardware constraints impose strict limits on circuit depth and coherence times, making deeper quantum networks computationally expensive and impractical [63, 64].
Thirdly, these qualities are especially advantageous for PINNs, where shallow network architectures are favored to optimize the balance between computational efficiency and solution accuracy [65, 66, 67, 5].
Finally, a substantial benefit of such hybrid models is the integration of automatic differentiation within both classical and quantum networks. This capability, supported by contemporary software packages like PennyLane [68] and TensorFlow Quantum [69], enables the automatic computation of derivatives as variables are directly incorporated into the computational graph. This approach simplifies PINN implementation and facilitates the efficient handling of complex physics-informed loss functions and higher-order derivatives.
Our work is related to the approaches in [12, 13, 14, 15], but with several key differences. While HQPINN [12] focuses specifically on computational fluid dynamics in 3D Y-shaped mixers using a fixed quantum depth with an infused layer structure, the QPINN [13] integrates a dynamic quantum circuit combining Gaussian and non-Gaussian gates with classical computing methodologies within a PINN framework to solve optimal control problems. In addition, the model proposed in [14] investigates both purely quantum and hybrid PINNs, highlighting parameter efficiency but relying on limited PQC networks and addressing primarily simple PDEs. Recently proposed HQPINN [15], inspired by Fourier neural operator (FNO) [70], introduces a parallel hybrid architecture to separately model harmonic and non-harmonic features before combining them for output.
4 Quantum-Classical Physics-Informed Neural Networks (QCPINN)
Figure 1: Architecture of the quantum-classical physics-informed neural network (QCPINN) framework. The figure illustrates the hybrid structure: a classical preprocessor encodes the inputs, which are then fed into the QNN or classical neural network (NN). The QNN is implemented as either a CV-circuit or a DV-circuit with $k$ layers. A classical postprocessor then decodes the QNN or NN outputs. The integrated loss function concurrently minimizes PDE residuals and initial/boundary condition errors, ensuring the model adheres to physical constraints throughout training. Table 1 lists the various configurations of the CV and DV-circuits studied in this work.
Table 1: This work examines different configurations of the QCPINN along with the list of PDEs used for testing and generalization.
QNN Configurations PDEs
Type Options Testing Generalization
CV-circuit Measurement scheme Number, quadrature
Helmholtz,
Lid-driven cavity
Wave,
Klein-Gordon,
Convection-diffusion
Nonlinear operations Kerr, cubic phase, cross-kerr
Parameterization
Phase-free (magnitude),
Full (magnitude and phase)
(also refer to algorithm 1)
DV-circuit Embeddings Amplitude, angle
Topology
Alternate, cascade, cross-mesh, layered
(also refer to Fig. 2 and table 2)
Refer to caption Refer to caption
(a) alternate (b) cascade
(c) cross-mesh
(d) layered
Figure 2: DV-circuit topologies evaluated in this study, shown with five qubits and angle embedding as an example. Square boxes denote parameterized single-qubit rotation gates, and lines with dots indicate entangling operations (e.g., CNOT gates). The characteristics of these circuits are shown in table 2.
Fig. 1 illustrates a high-level overview of our proposed QCPINN. The preprocessor consists of a two-layer neural network architecture, ensuring that the overall framework is sufficiently shallow for efficient training while also being deep enough to capture essential transformations. Each layer is composed of 50 neurons. The first layer performs a linear transformation to modify the input dimension, followed by a $Tanh$ activation function to introduce non-linearity. The second layer subsequently maps the transformed inputs to the QNN layer (or classical NN layer used for comparison).
Likewise, the postprocessor layer replicates the preprocessor structure with appropriate dimensions, ensuring a symmetric transformation between classical and quantum representations. This design offers flexibility in managing quantum feature representations while ensuring the higher-order differentiability required for residual loss in the proposed QCPINN model.
Algorithm 1 CV-circuit QNN
1:Qumodes: num_qumodes, Layers: num_layers, Device: $d$, Cutoff: $c$, Input: Tensor $x$
2:Output: Tensor $y$
3:Initialize Parameters:
[…]
5:Quantum Device: Initialize $d$ with num_qumodes, cutoff $c$.
6:procedure QuantumCircuit(input)
7:  for $i=1$ to num_qumodes do
8:   Encode Inputs with Displacements $D$: $D(\text{input}[i],0)$ $\triangleright$ only real amplitudes
9:  end for
10:  for $l=1$ to num_layers do
11:   Apply Interferometer($\psi_{1}[l]$ , $\phi_{1}[l]$)
12:   Apply Squeezing($r[l],\phi_{\text{s}}[l]$) $\triangleright$ Phase-free: $\phi_{\text{s}}[l]=0.0$
13:   Apply Interferometer($\psi_{2}[l]$ , $\phi_{2}[l]$)
14:   Apply Displacement($\alpha[l],\phi_{\text{d}}[l]$) $\triangleright$ Phase-free: $\phi_{\text{d}}[l]=0.0$
15:   Apply Nonlinearity $\Phi$ $\triangleright$ Options: Kerr, Cubic Phase, Cross-Kerr
16:  end for
17:  Measure state for each wire $\triangleright$ Measurement options: $\langle\hat{q}_{i}\rangle$ (Quadrature) or $\langle\hat{n}_{i}\rangle$ (Number)
18:  return Measurements
19:end procedure
20:procedure Interferometer($\psi$ , $\phi_{\text{bs}}[l]$)
21:  for $l=1$ to num_qumodes do
22:   for $(q_{1},q_{2})$ in Pairs(num_qumodesdo
23:     if $(l+k)\%2\neq 1$ then
24:      Apply Beamsplitter $B(\psi[n],\phi_{\text{bs}}[n])$ $\triangleright$ Phase-free: $\phi_{\text{bs}}[n]=0.0$
25:     end if
26:   end for
27:  end for
28:  for $i=1$ to $\text{num\_qumodes}-1$ do
29:   Apply Rotation $R_{Z}(\psi[i])$
30:  end for
31:end procedure
32:procedure ForwardPass(Input $x$)
33:  $y\leftarrow\text{QuantumCircuit}(x)$
34:  return y
35:end procedure
4.1 CV-circuit QCPINN
The CV-circuit QCPINN consists of three primary components: a preprocessor network, a CV-circuit quantum network (the QNN layer of Fig. 1), and a postprocessor network. The quantum network is adapted from [17] and is described in algorithm 1. Each layer of the quantum network executes a sequence of quantum operations.
Table 1 presents various configurations of our CV-circuit studied to identify a suitable parameterization scheme, nonlinear operation, and measurement approach. For measurement schemes, we evaluated both number and quadrature. For parameterization, we examined phase-free using only magnitude parameters and full parameterization using both magnitude and phase parameters. For nonlinear operations, we explored Kerr nonlinearity, cubic phase gates, and cross-Kerr interactions. Additionally, the quantum state is measured, typically in either the quadrature basis $\langle\hat{q}_{i}\rangle$ or the Fock number basis $\langle\hat{n}_{i}\rangle$, to yield classical outputs for the postprocessing layers. Please refer to section 2.2 for additional details.
Structuring the network as described in algorithm 1 results in a circuit depth of $(2n+3)L$, while the number of trainable quantum parameters increases to $\mathcal{O}(n^{2}L)$, where $n$ represents the number of qumodes and $L$ denotes the number of layers.
4.2 DV-circuit QCPINN
Table 2: Characteristics of the topologies used in DV-circuits in Fig. 2. $n$ is the number of qubits, and $L$ is the number of layers.
Topology Circuit Depth Circuit Connectivity Number of Parameters Number of two-qubit gates
alternate $6L$ Nearest-neighbor 4(n1)L4(n-1)L (n1)L(n-1)L
cascade $(n+2)L$ Ring topology $3nL$ nLnL
cross-mesh $(n^{2}-n+4)L$ All-to-all $(n^{2}+3n)L$ $(n^{2}-n)L$
layered $6L$ Nearest-neighbor 4(n1)L4(n-1)L (n1)L(n-1)L
The DV-circuit QCPINN consists of three primary components: a preprocessor network, a DV-circuit quantum network (the QNN layer of Fig. 1), and a postprocessor network. The quantum network is a custom DV-circuit, which executes quantum operations on the preprocessed data.
Table 1 presents various configurations of our DV-circuit studied to identify a suitable embedding scheme and topology. Fig. 2 illustrates four proposed ansätze named after their topology: alternate, cascade, cross-mesh, and layered. These circuits are adapted from literature, including  [71, 72, 73, 62]. Table 2 presents features of the DV-circuits across the key descriptors, circuit depth, connectivity, number of parameters, and the number of two-qubit gates. These circuits agree with the affine transformation discussed in section 2 and adhere to the hardware-efficient ansatz (HEA) [74] defined as:
$U(\psi)\;=\;\prod_{k}U_{k}(\psi_{k})\,W_{k},$ (3)
where, each layer $k$ alternates between single-qubit parameterized rotations, $U_{k}(\psi_{k})$, and an entangling operation, $W_{k}$.
The cross-mesh ansatz implements an all-to-all connectivity scheme, facilitating extensive interaction between qubits, which enhances its ability to explore the Hilbert space. The circuit depth scales as $\mathcal{O}(n^{2}-n+4)L$, while the number of trainable parameters is $(n^{2}+3n)L$. The use of global entanglement fosters strong correlations between qubits; however, it results in greater circuit depth, resource demands, and training difficulties [20]. This circuit’s total number of two-qubit gates is also significantly higher than in other ansätze.
The cascade ansatz balances expressivity and hardware efficiency by employing a ring topology for qubit connectivity. This design facilitates efficient entanglement while keeping the circuit depth manageable at $(n+2)L$. The circuit contains $3nL$ trainable parameters, making it more efficient than cross-mesh while maintaining reasonable expressivity. Incorporating controlled rotation gates ($CR_{X}$) improves entanglement compared to CNOT gates and increases the circuit’s capacity to explore the Hilbert space. Furthermore, the requirement for two-qubit gates is fewer than that of cross-mesh, thus lowering the overall computational cost.
The layered ansatz features a structured design that includes alternating rotation and entangling layers. Each qubit undergoes parameterized single-qubit rotations, typically utilizing $R_{X}$ and $R_{Z}$ gates. Entanglement is achieved through nearest-neighbor CNOT gates, ensuring compatibility with hardware constraints while maintaining efficient connectivity. Compared to the cross-mesh’s all-to-all connectivity or the cascade’s ring topology, the layered architecture represents a compromise between expressivity and hardware efficiency.
The alternate ansatz uses an alternating-layer structure featuring nearest-neighbor CNOT gates, ensuring effective entanglement while maintaining a manageable circuit depth of $6L$. The total number of trainable parameters in this circuit 4(n1)L4(n-1)L aligns with the layered ansatz. The entanglement structure determined by the alternating CNOT layers offers the highest level of entangling capability among the circuits studied. Consequently, the cross-mesh ansatz provides the largest parameter count and highest expressivity, whereas the alternate circuit excels in entangling capabilities.
Compared to the QNN models discussed above, we can see that the classical feed-forward neural networks (NNs) exhibit a more flexible scaling pattern, with trainable parameters $\mathcal{O}(h^{2}L)$, where $h$ denotes the number of hidden neurons (assuming a fixed width $h$ in all hidden layers) with independent selection of $h$. Therefore, while classical models depend on deep architectures and larger parameter counts, quantum networks achieve expressivity by utilizing intrinsic quantum correlations and exponential scaling of the Hilbert space.
5 Implementation and Training Details
We used PyTorch’s automatic differentiation to compute the gradients for both classical and quantum parameters. The training was carried out with a batch size of 64 and a maximum of 20,000 epochs in all case studies, except in the cavity problem, where we used 12,500 maximum iterations. The Adam optimizer was used to ensure stable convergence. These settings were uniformly applied across all study cases to provide a consistent basis for comparison. For the cavity problem, initial training data was selected using the Sobol sequence, and then random sampling with a batch size of 128 was applied to this Sobol-generated dataset during training. For all other cases, only random sampling with a batch size of 128 was used throughout the entire process.
While other training configurations, such as increasing the number of iterations or adjusting the batch size, could yield better performance, the goal is to evaluate the models under identical conditions. Given the near-infinite hyperparameter space, this controlled evaluation ensures a fair assessment of the relative strengths and limitations of each approach. It is also important to acknowledge that the quantum experiments conducted in this study were performed using emulator-based simulations rather than actual quantum hardware. These simulations inherently operate at significantly reduced speeds compared to their classical counterparts due to the substantial computational overhead associated with emulating quantum circuits on classical machines. Future implementations on actual quantum hardware would eliminate this simulation overhead, potentially revealing additional performance characteristics not captured in the current study.
In our initial implementation, we observed that the experiments on the CPU were faster than those on the GPU. This is likely due to our study’s relatively small batch size, circuit complexity, and limited number of qubits/qumodes: five qubits for DV-circuit models and two qumodes for CV-based models. Moreover, the quantum backends employed in our experiments are not optimized for GPU efficiency. Considering these limitations, we have chosen to conduct our experiments on a CPU system with 6148 CPUs (each operating at 2.40 GHz) and 192 GB of RAM. Further details on training configurations and implementation strategies are provided in the following subsections.
5.1 CV-circuit QCPINN
We used the PennyLane framework [68] with the Strawberry Fields Fock backend [75] to develop and train our CV-circuit QCPINN model. Each CV-circuit layer was configured for two qumodes with a cutoff dimension of 20. Quantum parameters were initialized using controlled random distributions, with small values for active parameters (e.g., $r$, $d$) to ensure numerical stability. Classical PINN’s layers were initialized using Xavier initialization. During training, either position quadrature (X) or number operator expectation measurements are performed on each qumode to extract continuous-valued features from the quantum state.
The CV-circuit operations are carefully managed to ensure numerical stability. For instance, squeezing operations are highly sensitive to large values, as excessive squeezing amplifies quadrature fluctuations, which increases photon number variance and truncation errors in Fock-space representations [76]. Similarly, displacement operations are constrained to prevent quantum states from exceeding computational cutoffs, as this can introduce simulation errors [77]. Likewise, the nonlinear operations such as Kerr gate can cause phase shifts that scale quadratically with photon numbers, making them highly sensitive to large values [78].
Therefore, we used a learning rate of $1.0\times 10^{-4}$ to address the model’s sensitivity to gradient updates. A ReduceLROnPlateau scheduler was implemented with a patience of 20 epochs, decreasing the learning rate by a factor of 0.5 when no improvement was observed, down to a minimum of $1.0\times 10^{-6}$. Additionally, we applied gradient clipping to prevent exploding gradients, restricting their norm to a maximum of 0.1. As shown in algorithm 1, we also scaled the training parameters by small factors to further stabilize the CV transformations.
While these settings can alleviate the instability and prevent divergence in quantum state evolution, they can also introduce the vanishing gradient problem. This issue is well-documented in classical neural network literature and occurs when excessively small initialization values diminish gradients during backpropagation, slowing or even halting the learning process [79].
5.2 DV-circuit QCPINN
We used Adam optimizer with a learning rate of 0.005 to train the DV-circuit QCPINN. A learning rate scheduler, ReduceLROnPlateau, is employed to adjust the learning rate when the loss plateaus dynamically. Specifically, the scheduler reduces the learning rate by a factor of 0.9 if no improvement is observed for 1000 consecutive epochs, promoting stability and convergence. Furthermore, the training function applies gradient clipping to prevent exploding gradients, clipping them according to their norm to ensure they do not exceed a unity value. During training, Pauli-Z expectation measurements are performed on each qubit to extract features from the quantum state.
5.3 Classical PINN Model
The classical baseline model consists of three main components: a preprocessing network, a neural network NN, and a postprocessing network (as shown in figure 1). The NN differs between two configurations: (a) In model-1, it includes two fully connected layers, each containing 50 neurons with a Tanh activation function between layers, and (b) In model-2, the NN is entirely omitted, meaning only the preprocessing and postprocessing networks are kept, simplifying the model’s structure.
These two models represent different levels of complexity, with model-1 having more layers and, consequently, a higher number of trainable parameters than model-2. We use these classical models to evaluate and compare the performance of the proposed hybrid models in figure 1. Training settings, including the optimizer and batch size, were kept consistent with the QCPINN models to allow for a fair comparison between classical and quantum-enhanced approaches. The learning rate is $0.005$.
5.4 PDEs
We implemented five distinct PDEs, which include the Helmholtz, the time-dependent 2D lid-driven cavity, the 1D wave, the Klein-Gordon, and a convection-diffusion problems. The mathematical formulation, boundary conditions, and corresponding loss function designs for each PDE are detailed in A.
Table 3: The CV-circuit QCPINN models the performance in solving the Helmholtz and lid-driven cavity equations. There are 12 combinations in total, as in Table 1. Among these results of the model with a quadrature measurement scheme, a Kerr nonlinear function, and phase-free parameterization is reported here. The remaining 11 combinations are not mentioned as the results are more or less similar. Training convergence is shown in Fig. 3.
PDE # Parameters   Relative $\mathbf{L_{2}}$ error (%) Final loss
Helmholtz 469 $u$ $1.00e+02$ 5077.79
$f$ $1.00e+02$
cavity 621 $u$ $9.345e+01$ 0.102
$v$ $9.979e+01$
pp $8.620e+01$
Refer to caption Refer to caption
(a) Helmholtz (b) cavity
Figure 3: The training convergence behavior of the CV-circuit QCPINN model when solving (a) Helmholtz and (b) lid-driven cavity. The terms “loss_r” and “loss_bc” refer to residual and boundary losses. The “loss_bc” shows a decreasing trend, indicating improvements in satisfying the boundary conditions; however, the overall decrease is below expectation. The “loss_r” exhibits significant fluctuations and does not display a noticeable reduction, suggesting difficulties in minimizing the residual error of the governing equation.
6 Results
We began testing our CV and DV-circuit QCPINN using two benchmark problems: Helmholtz and the lid-driven cavity.
6.1 CV-circuit QCPINN
At the beginning of our study, we implemented CV-circuit QCPINN models using 12 different combinations. This included two measurement schemes, three nonlinear functions, and two parameterization schemes (see Table 1). The results indicate that the behavior of these models is quite similar. Table 3 and Fig. 3 present the results and training convergence for the model that employs the quadrature measurement scheme, Kerr nonlinear function, and phase-free parameterization.
6.1.1 Helmholtz Equation.
For the Helmholtz Eq. 4, we used 469 trainable parameters. The results, in Table 3, show a high relative $L_{2}$ error for both the $u$ and $f$ variables, along with a significant final loss. The training convergence pattern in Fig. 3(a) highlights some challenges. For example, the boundary loss, “loss_bc”, shows a steady decrease over 10,000 iterations, suggesting that the model is gradually improving in meeting the boundary conditions. However, the residual loss, “loss_r”, remains consistently high, with substantial fluctuations throughout the training process. This indicates that the model was struggling to satisfy the differential equation itself.
6.1.2 Time-dependent 2D Lid-driven Cavity Equation.
For the more complex lid-driven cavity Eq. 5, the CV-circuit QCPINN model was implemented with 621 trainable parameters. The result, in Table 3, shows that the relative $L_{2}$ errors for both velocity and pressure variables remain high. However, the final loss value of 0.102 is significantly lower than that observed for the Helmholtz equation. Fig. 3(b) shows the training convergence, where the boundary loss exhibits some improvement, although it still experiences considerable fluctuations. The residual loss displays even greater oscillations compared to the Helmholtz case, indicating a greater challenge in accurately capturing the complex fluid dynamics governed by the Navier-Stokes equations.
This behavior is primarily attributed to gradient decay issues intrinsic to our CV-circuit implementation. The vanishing gradient problem [79] manifests more severely due to the additional constraints imposed by quantum operations. Small initialization values, while necessary to prevent divergence in quantum state evolution, significantly diminish gradients during backpropagation.
We examined various techniques, such as Xavier initialization for learnable parameters, gradient clipping, and skip connections, to mitigate the problem of gradient decay, but these attempts proved ineffective. Exploring alternative optimization strategies beyond standard gradient descent may be beneficial in addressing this issue, though this remains beyond the scope of this work.
6.2 DV-circuit QCPINN
The DV-circuit QCPINNs demonstrate stable performance compared to the CV-circuit across all the PDEs considered in this study. Thus, we perform a detailed analysis of the DV-circuit implementation throughout the remainder of this paper.
Our investigation began by using various DV-circuit configurations to solve two fundamental PDEs: Helmholtz and the lid-driven cavity. We systematically evaluated the performance across combinations of two distinct embedding approaches and four different circuit topologies, comparing these quantum implementations against classical models.
Table 4: Comparison of the performance of our DV-circuit QCPINN models against classical PINN approaches in solving the Helmholtz equation. The comparison metrics include the number of trainable parameters required by each model, the relative $L_{2}$ error in %, and the final training loss value achieved. $u$ and $f$ are the PDE solution and the force term.
Embedding Topology # Parameters   Relative $\mathbf{L_{2}}$ error (%) Final loss
DV-circuit amplitude alternate 781 $u$ $2.14e+01$ $55.360$
$f$ $1.21e+01$
cascade 771 $u$ $1.99e+01$ $6.372$
$f$ $4.97e+00$
cross-mesh 836 $u$ $1.77e+01$ $6.872$
$f$ $3.76e+00$
layered 781 $u$ $4.05e+01$ $39.441$
$f$ $7.54e+00$
angle
alternate 781 $u$ $9.15e+00$ $11.456$
$f$ $4.15e+00$
cascade 771 $u$ $4.12e+00$ $3.301$
$f$ $1.64e+00$
cross-mesh 836 $u$ $1.18e+01$ $5.822$
$f$ $3.76e+00$
layered 781 $u$ $2.35e+01$ $11.512$
$f$ $5.73e+00$
Classical
PINN
model-1 7851 $u$ $3.18e+01$ $20.331$
$f$ $6.40e+00$
model-2 2751 $u$ $6.72e+00$ $4.771$
$f$ $2.83e+00$
Figure 4: Training convergence history of our DV-based QCPINN models compared to best classical PINN model when solving the Helmholtz equation. The plots track the performance over 20,000 epochs, showing how QCPINN performs with two different embedding methods across four circuit topologies. For comparison, we include the classical PINN model-2, which outperforms model-1 in this case.
Figure 5: Comparison of the exact solution and the model predictions for the Helmholtz equation: the best classical PINN results (model-2), and our DV-circuit QCPINN using angle embedding (which outperforms amplitude embedding) with cascade topology (which outperforms the other three topologies). The first row shows the PDE solution $u$, while the second row shows the force field ($f$).
6.2.1 Helmholtz Equation.
Table 4 presents the performance of the DV-circuit QCPINN models in solving the Helmholtz Eq. 4 and comparing the results with classical models. The results show that the cascade topology consistently outperformed other topologies, achieving the lowest final training loss and low relative $L_{2}$ testing errors. In contrast, the layered and alternate circuit topologies show higher final training losses, especially the layered topology with amplitude embedding.
Fig. 4 compares the training loss convergence of DV-circuit QCPINN models to the classical PINN. The results indicate that DV-based quantum models display controlled convergence with relatively stable training dynamics. Moreover, the angle embedded cascade topology demonstrates the most competitive performance, achieving a loss level close to the best classical model while utilizing approximately 72% fewer parameters.
Fig. 5 visualizes the exact and predicted solutions for the Helmholtz equation. It also shows the prediction errors of both the best classical model and the angle-cascade QCPINN. While the classical model captures the overall structure, it exhibits noticeable discrepancies, as shown in the absolute error maps, which reveal significant deviations near the boundaries and at regions of sharp local gradients. In contrast, the predictions from the angle-cascade QCPINN model closely align with the exact solutions while using fewer parameters than the classical approach.
Table 5: Comparison of the performance of our DV-circuit QCPINN models against classical PINN approaches in solving the lid-driven cavity equation. The comparison metrics include the number of trainable parameters required by each model, the relative $L_{2}$ error in %, and the final training loss value achieved. $u$, $v$, and $p$ are the velocity components and the pressure values, respectively.
Embedding Topology # Parameters   Relative $\mathbf{L_{2}}$ error (%) Final loss
DV
amplitude
alternate 933 $u$ $6.441e+01$ 0.127
$v$ $8.439e+01$
pp $6.102e+01$
cascade 928 $u$ $6.670e+01$ 0.123
$v$ $8.409e+01$
pp $6.365e+01$
cross-mesh 988 $u$ $7.073e+01$ 0.109
$v$ $4.893e+01$
pp $7.096e+01$
layered 933 $u$ $5.690e+01$ 0.121
$v$ $9.476e+01$
pp $7.469e+01$
angle
alternate 933 $u$ $4.041e+01$ 0.109
$v$ $4.683e+01$
pp $4.746e+01$
cascade 928 $u$ $2.745e+01$ 0.083
$v$ $2.081e+01$
pp $4.791e+01$
cross-mesh 988 $u$ $2.035e+01$ 0.084
v $2.071e+01$
p $4.011e+01$
layered 933 $u$ $4.572e+01$ 0.129
v $6.628e+01$
p $7.885e+01$
Classical
PINN
model-1 8003 $u$ $2.103e+01$ 0.087
$v$ $3.027e+01$
pp $2.317e+01$
model-2 2903 $u$ $3.830e+01$ 0.073
$v$ $2.924e+01$
pp $4.106e+01$
Figure 6: Training convergence history of our DV-based QCPINN models compared to the best classical PINN model (model-2) when solving the lid-driven cavity equation. The plots track the performance over 12500 epochs, showing how QCPINN performs with two different embedding methods across four circuit topologies.
Figure 7: Comparison of the reference FEM solution and the model predictions for the lid-driven cavity equation: the classical PINN model-2 (which outperforms model-1), and our DV-circuit QCPINN using angle embedding (which outperforms amplitude embedding) with cascade topology (which outperforms the other three topologies). The first and second rows show the velocity fields ($u$ and $v$), while the third row shows the pressure field ($p$).
6.2.2 Time-dependent 2D Lid-driven Cavity Equation.
Table 5 presents the performance of the QCPINN methods for solving the lid-driven cavity Eq. 5. Among the DV-circuit QCPINN models, cascade achieves the best overall performance, demonstrating good convergence properties with the lowest final training loss. The models with angle embedding and cross-mesh topology also perform well, achieving the lowest final training loss and highlighting their effectiveness in this setting. Specifically, the angle-cascade quantum model achieves comparable performance to the best classical model while requiring only 12% of the parameters.
Fig. 6 compares the training loss convergence of various DV-circuit QCPINN models with a classical PINN while solving the lid-driven cavity equation. The loss curves indicate that all DV-based models achieve stable convergence with minimal oscillations. Similar to Fig. 6, the angle embedding with the cascade topology circuit in Fig. 4 closely matches the loss level of the best classical model.
Fig. 7 visualizes the results obtained using the finite element method solution alongside the predicted solutions for the lid-driven cavity Eq. 5. It also shows the prediction errors for both the best classical model and the angle-cascade QCPINN. The classical PINN model and the angle-cascade QCPINN effectively capture the main flow features. The absolute error distributions (third and fifth rows) reveal distinct patterns in both velocity components and pressure fields. Both models exhibit higher errors near the top moving lid for velocity components, where the flow experiences the most significant shear.
6.2.3 Generalizing Our Solution.
Table 6: Comparison of our DV-circuit QCPINN using angle embedding (which outperforms amplitude embedding) with cascade topology (which outperforms the other three topologies) with the two classical PINN models for solving wave, Klein-Gordon, and convection-diffusion equations. The comparison metrics include the number of trainable parameters required by each model, the relative $L_{2}$ error in %, and the final training loss value achieved. $u$ and $f$ are the reference solutions and the source term, respectively.
PDEs Method
# Parameters
Relative $\mathbf{L_{2}}$ error (%)
Final loss
wave Eq. 6 angle-cascade 771 $u$ $1.02e+01$ $0.083$
Classical
PINN
model-1 $7851$ $u$ $2.06e+01$ $0.128$
model-2 2751 $u$ $8.68e+00$ $0.027$
Klein-Gordon
Eq. 7
angle-cascade 771 $u$ $1.17e+01$ $2.727$
$f$ $3.17e+00$
Classical
PINN
model-1 7851 $u$ $1.75e+01$ $0.718$
$f$ $3.35e+00$
model-2 2751 $u$ $1.76e+01$ $0.358$
$f$ $3.78e+00$
Convection-diffusion
Eq. 8
angle-cascade 821 $u$ $1.20e+01$ $0.008$
$f$ $4.38e+00$
Classical
PINN
model-1 7901 $u$ $2.00e+01$ $0.014$
$f$ $3.11e+00$
model-2 2801 $u$ $2.06e+01$ $0.1$
$f$ $1.31e+01$
Refer to caption Refer to caption Refer to caption
(a) wave (b) Klein-Gordon (c) Convection-diffusion
Figure 8: Training loss history for three PDEs: (a) wave, (b) Klein-Gordon, and (c) convection-diffusion. Each plot compares the convergence behavior of our angle-cascade QCPINN model (which outperforms the other embedding and topologies) against the best-performing classical PINN model (as reported in table LABEL:tab:more_cases) over 20,000 training epochs.
(a) wave
(b) Klein-Gordon
(c) convection-diffusion
Figure 9: Comparison of the reference solution and the model predictions for three PDEs: (a) wave, (b) Klein-Gordon, and (c) convection-diffusion. For each equation, we present results from both the best-performing classical PINN model and angle-cascade DV-circuit QCPINN implementation. For each equation, the top row displays the velocity field ($u$), while the bottom row (if exists) shows the force field ($f$).
Our experiments with the Helmholtz and lid-driven cavity problems, outlined in sections 6.2.1 and 6.2.2, revealed that the angle-cascade variant of the DV-circuit QCPINN architecture consistently outperforms other configurations. To validate the robustness of the angle-cascade QCPINN architecture across diverse physical systems, we expanded our evaluation to three additional canonical PDEs: the wave equation 6, the Klein-Gordon equation 7, and the convection-diffusion equation 8. This expansion allows us to assess whether the advantages of our quantum-enhanced architecture persist across problems with fundamentally different physical characteristics and computational challenges.
Table LABEL:tab:more_cases presents the performance of the angle-cascade DV-circuit QCPINN model for solving these three additional PDEs and comparing the results with the classical models. Our angle-cascade QCPINN model maintains its parameter efficiency advantage while achieving competitive or better accuracy compared to classical PINNs. For the wave equation, our model requires only a 9.8% parameters while achieving a 50.5% reduction in relative $L_{2}$ error. The Klein-Gordon equation presents additional complexity due to its coupled system nature. Although our model exhibits a slightly higher final training loss, it maintains comparable relative $L_{2}$ errors while using only 9.8% of the parameters required by model-1. Notably, for the convection-diffusion equation, angle-cascade QCPINN model with 821 parameters (compared to 7,901 in model-1) achieves a 43% reduction in relative $L_{2}$ error for the primary solution variable and demonstrates improved training convergence.
Fig. 8 compares the convergence behavior of the angle-cascade DV-circuit QCPINN model to the classical PINN model. For the wave equation, Fig. 8(a) shows that the angle-cascade QCPINN (blue line) demonstrates significantly faster initial convergence, achieving lower loss values within the first 5,000 epochs compared to the classical approach (pink line). Likewise, for the Klein-Gordon equation, Fig. 8(b) shows that angle-cascade QCPINN maintains comparable convergence stability despite the equation’s coupled nature, with both models exhibiting similar loss trajectories after the initial training phase. Notably, for the convection-diffusion equation, Fig. 8(c) shows that our model consistently maintains a monotonic loss reduction throughout training, reaching a final loss approximately one order of magnitude lower than that of the classical approach.
Fig. 9 visualizes both reference and predicted solutions. For the wave equation, Fig. 9(a) shows that both the classical PINN and QCPINN capture the oscillatory pattern, but the error distribution of QCPINN (top-right plot) displays noticeably lower error magnitudes across the domain. The Klein-Gordon equation, Fig. 9(b), shows our model’s ability to accurately represent coupled solutions, with error distributions that are comparable to or better than those of the classical model. The convection-diffusion equation, Fig. 9(c), shows our model’s ability to precisely capture the sharp localized features characteristic of convection-dominated flows. The error visualization confirms significantly reduced prediction errors across solution components ($u$ and $f$), with error magnitudes approximately 232-3 times lower than those of classical PINNs.
7 Conclusion
We found that QCPINN models achieve comparable solution accuracy to classical PINNs while requiring a fraction of trainable parameters – as little as 10% of those needed by their classical counterparts under about the same settings. The experimental evidence supports our hypothesis that hybrid quantum-classical approaches can maintain solution quality while dramatically improving parameter efficiency. Our systematic assessment of various QCPINN configurations across multiple PDEs consistently revealed that DV-circuit implementations with angle embedding and cascade topology offer the best performance. Notably, for the convection-diffusion equation, our angle-cascade QCPINN achieved around 40% reduction in relative $L_{2}$ error compared to classical models.
Our experimental results reveal that CV-circuits consistently show training instabilities and underperform compared to DV-circuits. Standard gradient-based optimization methods, like Adam, face significant challenges with CV-based quantum circuit models due to their sensitivity to parameter initialization and gradient dynamics. The vanishing gradient problem is worsened in CV circuits, where small initial parameter values are needed for stability, but these values also limit gradient magnitudes during backpropagation, creating a trade-off between stability and trainability.
The primary strength of this work lies in its comprehensive comparative study, which systematically evaluates two quantum circuit architectures: DV and CV, each assessed with various configurations. The CV circuits comprise two measurement schemes, three nonlinear operations, and two parameterization schemes, while the DV circuits consist of two embedding schemes and four circuit topologies, evaluated against two benchmark PDEs and generalizability against three additional benchmark PDEs. The open-source implementation further enhances reproducibility and enables community engagement. However, we acknowledge several limitations. While parameter efficiency was significantly improved, overall solution accuracy remained comparable to classical PINNs rather than surpassing them, except for the convection-diffusion problem. Additionally, the highly non-convex PINN loss landscape still required careful weighting for the loss terms, which were determined through extensive empirical studies rather than exploring quantum circuit advantage to enhance this problem. Contrary to theoretical expectations, CV-circuit implementations exhibited training instabilities and underperformed relative to DV-circuit variants for the cases considered in this study.
Despite these limitations, our findings remain robust when examined through multiple performance metrics and across diverse PDE types. The consistent advantage of angle-cascade configurations in achieving lower relative $L_{2}$ errors and final training losses, combined with their significant parameter efficiency, represents a meaningful advancement in physics-informed machine learning.
Future research should explore more complex PDEs that require larger parameter spaces. By applying PINNs and QPINNs to these scenarios, a comparative analysis could highlight the benefits of QCPINNs and reveal advantages not evident in simpler domains or that PINNs cannot address. This broader investigation could enhance our understanding of when quantum-enhanced neural networks significantly improve solutions for complex PDEs. Additionally, studies should focus on advanced quantum circuit designs tailored to PDE characteristics and on developing optimization strategies to improve training stability, especially for CV-circuit quantum methods.
Acknowledgment
We thank the National Center for High-Performance Computing of Turkiye (UHeM) for providing computing resources under grant number 5010662021.
Keywords Quantum Computing  $\cdot$ Neural networks  $\cdot$ Physics-informed neural network  $\cdot$ Partial differential equation  $\cdot$ PINN  $\cdot$ PDE
Appendix A PDEs
A.1 Helmholtz Equation
The Helmholtz equation represents a fundamental PDE in mathematical physics, arising as the time-independent form of the wave equation. We consider a two-dimensional Helmholtz equation of the form:
$\begin{split}\Delta u(x,y)+k^{2}u(x,y)&=f(x,y)\qquad(x,y)\in\Omega\\ u(x,y)&=h(x,y)\qquad(x,y)\in\Gamma_{0}\end{split}$ (4)
where $\Delta$ is the Laplacian operator defined as $\Delta u=\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}$, $k$ is the wavenumber (related to the frequency of oscillation), $f(x,y)$ is the forcing term (source function), $h(x,y)$ specifies the Dirichlet boundary conditions. In this study, we choose an exact solution of the form:
$u(x,y)=\sin(a_{1}\pi x)\sin(a_{2}\pi y)$
on the square domain $\Omega=[-1,1]\times[-1,1]$ with Wavenumber $k=1$, first mode number $a_{1}=1$,and the second mode number $a_{2}=4$. This choice of solution leads to a corresponding source term:
$f(x,y)=u(x,y)[k^{2}-(a_{1}\pi)^{2}-(a_{2}\pi)^{2}]$
To solve this problem using QCPINN, we construct a loss function that incorporates both the physical governing equation and the boundary conditions:
$\mathcal{L}(\theta)$ $=\underset{\theta}{\min}\big{(}\lambda_{1}\|\mathcal{L}_{\text{phy}}(\theta)\|_{\Omega}+\lambda_{2}\|\mathcal{L}_{\text{bc}}(\theta)\|_{\Gamma_{0}}\big{)}$
$=\underset{\theta}{\min}\big{(}\lambda_{1}\text{MSE}(\|\underbrace{u_{\theta_{xx}}(x,y)+u_{\theta_{yy}}(x,y)+\alpha u_{\theta}(x,y)}_{\text{PDE residual}}\|_{\Omega})+\text{MSE}(\lambda_{2}\|\underbrace{u_{\theta}(x,y)-u(x,y)}_{\text{Boundary condition}}\|_{\Gamma_{0}})\big{)}$
where $\alpha=(a_{1}\pi)^{2}+(a_{2}\pi)^{2}$ combines the mode numbers into a single parameter. The chosen loss weights are $\lambda_{1}=1.0$ and $\lambda_{2}=10.0$.
A.2 Time-dependent 2D lid-driven cavity Problem
The lid-driven cavity flow is a classical CFD benchmark problem that is a fundamental test case for many numerical methods where flow is governed by the unsteady, incompressible Navier-Stokes equations. The momentum equation, continuity equation, initial condition, no-slip boundary on walls, and moving lid condition are defined, respectively, as :
$\begin{split}\rho\left(\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}\right)&=-\nabla p+\mu\nabla^{2}\mathbf{u}\\ \nabla\cdot\mathbf{u}&=0\\ \mathbf{u}(0,\mathbf{x})&=0\ \\ \mathbf{u}(t,\mathbf{x}_{0})&=0\qquad\mathbf{x}\in\Gamma_{0}\\ \mathbf{u}(t,\mathbf{x}_{l})&=1\qquad\mathbf{x}\in\Gamma_{1}\end{split}$ (5)
The computational domain is $\Omega=(0,1)\times(0,1)$, the spatial discretization, $(N_{x},N_{y})=(100,100)$ is uniform grid, with temporal domain $t\in[0,10]$ seconds and $\Delta t=0.01s$ with density $\rho=1056$ kg/m³, viscosity $\mu=1/\mathrm{Re}=0.01$ kg/(m·s), where Re is the Reynolds number. $\Gamma_{1}$ is the top boundary (moving lid) with tangential velocity $U=1$ m/s, and $\Gamma_{0}$ is the remaining three sides with no-slip condition ($\mathbf{u}=0$) where $\mathbf{u}=(u,v)$.
For validation purposes, we compare the neural network approximation with results obtained using the finite volume method. The PINN approach employs a composite loss function:
$\mathcal{L}(\theta)$ $=\underset{\theta}{\min}\big{[}\lambda_{1}\|\underbrace{\mathcal{L}_{\text{phy}}(\theta)}_{\text{PDE residual}}\|_{\Omega}+\lambda_{2}\|\underbrace{\mathcal{L}_{\text{up}(\theta)}+\mathcal{L}_{\text{bc}_{1}}(\theta)}_{\text{Boundary conditions}}\|_{\Gamma_{1}\cup\Gamma_{0}}+\lambda_{3}\|\underbrace{\mathcal{L}_{\text{u0}}(\theta)}_{\text{Initial conditions}}\|_{\Omega}\big{]}$
where, the physics-informed loss component, $\mathcal{L}_{\mathrm{phy}}(\theta)$ consists of three terms:
$\mathcal{L}_{\mathrm{phy}}(\theta)=\mathcal{L}_{r_{u}}+\mathcal{L}_{r_{v}}+\mathcal{L}_{r_{c}}$
such that,
$\mathcal{L}_{r_{u}}(\theta)$ $=\text{MSE}\left[(u_{{\theta}_{t}}+u_{\theta}u_{{\theta}_{x}}+v_{\theta}u_{{\theta}_{y}})+\frac{1.0}{\rho}p_{{\theta}_{x}}-\mu(u_{{\theta}_{xx}}+u_{{\theta}_{yy}})\right]$
$\mathcal{L}_{r_{v}}(\theta)$ $=\text{MSE}\left[(v_{{\theta}_{t}}+u_{\theta}v_{{\theta}_{x}}+v_{\theta}v_{{\theta}_{y}})+\frac{1.0}{\rho}p_{{\theta}_{y}}-\mu(v_{{\theta}_{xx}}+v_{{\theta}_{yy}})\right]$
$\mathcal{L}_{r_{c}}(\theta)$ $=\text{MSE}\left[u_{{\theta}_{x}}+v_{{\theta}_{y}}\right]$
The boundary and initial conditions are enforced through the following:
$\mathcal{L}_{\text{up}}$ $=\text{MSE}\left[(1.0-\hat{u})+\hat{v}\right$]
$\mathcal{L}_{\text{bc}_{1}}$ $=\mathcal{L}_{\text{bottom, right, left}}=\text{MSE}\left[\hat{u}+\hat{v}\right$]
$\mathcal{L}_{\text{u}_{0}}$ $=\text{MSE}\left[\hat{u}+\hat{v}+\hat{p}\right$]
where, $\mathcal{L}_{\mathrm{up}}$ is the moving lid, $\mathcal{L}_{\mathrm{bc1}}$ is the no-slip walls, $\mathcal{L}_{\mathrm{u_{0}}}$ is the initial conditions. The loss weights are empirically chosen as $\lambda_{1}=0.1$, $\lambda_{2}=2.0$, $\lambda_{2}=2.0$, $\lambda_{3}=4.0$ for $\mathcal{L}_{\mathrm{phy}}$, $\mathcal{L}_{\mathrm{up}}$ , $\mathcal{L}_{\mathrm{bc1}}$ and $\mathcal{L}_{\mathrm{u_{0}}}$ respectively.
A.3 1D Wave Equation
The wave equation is a fundamental second-order hyperbolic partial differential equation that models various physical phenomena. In its one-dimensional form, it describes the evolution of a disturbance along a single spatial dimension over time. The general form of the time-dependent 1D wave equation is:
$\begin{split}u_{tt}(t,x)-c^{2}u_{xx}(t,x)&=0\qquad\qquad(t,x)\in\quad\\ u(t,x_{0})&=f_{1}(t,x)\qquad(t,x)\text{ on }\quad\Gamma_{0}\\ u(t,x_{1})&=f_{2}(t,x)\qquad(t,x)\text{ on }\quad\Gamma_{1}\\ u(0,x)&=g(t,x)\qquad(t,x)\in\quad\Omega\\ u_{t}(0,x)&=h(t,x)\qquad(t,x)\in\quad\partial\Omega\end{split}$ (6)
where $u(t,x)$ represents the wave amplitude at position $x$ and time $t$, and $c$ is the wave speed characterizing the medium’s properties. The subscripts denote partial derivatives: $u_{tt}$ is the second time derivative, and $u_{xx}$ is the second spatial derivative.
[…]
$u(t,x)=\sin(\pi x)\cos(c\pi t)+0.5\sin(2c\pi x)\cos(4c\pi t)$
This solution represents a superposition of two standing waves with different spatial and temporal frequencies. The problem is defined on the unit square domain $(t,x)\in[0,1]\times[0,1]$, leading to the specific boundary value problem:
$u_{tt}(t,x)-4u_{xx}(t,x)$ $=0\qquad(t,x)\in\Omega=[0,1]\times[0,1]$
$u(t,0)=u(t,1)$ $=0$
$u(0,x)$ $=\sin(\pi x)+0.5\sin(4\pi x)\quad x\in[0,1]$
$u_{t}(0,x)$ $=0$
To solve this problem using the proposed hybrid QCPINN, we construct a loss function that incorporates the physical constraints, boundary conditions, and initial conditions:
$\mathcal{L}(\theta)$ $=\underset{\theta}{\min}\big{[}\lambda_{1}\|\mathcal{L}_{\text{phy}}(\theta)\|_{\Omega}+\lambda_{2}\|\mathcal{L}_{\text{bc}}(\theta)\|_{\Gamma_{1}}+\lambda_{3}\|\mathcal{L}_{\text{ic}}(\theta)\|_{\Gamma_{0}}\big{]}$
$=\underset{\theta}{\min}\big{[}\lambda_{1}\text{MSE}(\|\underbrace{u_{\theta_{tt}}(t,x)-4u_{\theta_{xx}}(t,x)}_{\text{PDE residual}}\|_{\Omega})$
$+\lambda_{2}\text{MSE}(\|\underbrace{u_{\theta}(t,0)+u_{\theta}(t,1)+u_{\theta}(0,x)-\sin(\pi x)-0.5\sin(4\pi x)}_{\text{Boundary/initial conditions}}\|_{{\Gamma_{0}}\cap{\Gamma_{1}}})$
$+\lambda_{3}\text{MSE}(\|\underbrace{u_{\theta_{t}}(0,x)}_{\text{Initial velocity}}\|_{\partial\Omega})\big{]}$
The loss weights are empirically chosen as $\lambda_{1}=0.1$, $\lambda_{2}=10.0$ and $\lambda_{3}=0.1$.
A.4 Klein-Gordon Equation
The Klein-Gordon equation is a significant second-order hyperbolic PDE that emerges in numerous theoretical physics and applied mathematics areas. The equation represents a natural relativistic extension of the Schrödinger equation. In this study, we consider the one-dimensional nonlinear Klein-Gordon equation of the form:
$\begin{split}u_{tt}-\alpha u_{xx}+\beta u+\gamma u^{k}&=0\qquad(t,x)\in\Omega\\ u(t,x)&=g_{1}(t,x)\qquad(t,x)\text{ on }\Gamma_{0}\\ u_{t}(t,x)&=g_{2}(t,x)\qquad(t,x)\text{ on }\Gamma_{1}\\ u(0,x)&=h(t,x)\qquad(t,x)\in\partial\Omega\times[0,T]\end{split}$ (7)
where $\alpha$ is the wave speed coefficient, $\beta$ is the linear term coefficient, $\gamma$ is the nonlinear term coefficient, and $k$ is the nonlinearity power. For our numerical study, we set $\alpha=1$ , $\beta=0$ $\gamma=1$ and $k=3$. We choose an exact solution of the form:
$u(t,x)=x\cos(5\pi t)+(tx)^{3}$
This solution combines oscillatory behavior with polynomial growth, providing a challenging test case for our numerical method. The complete boundary value problem on the unit square domain becomes:
$u_{tt}(t,x)-u_{xx}(t,x)+u^{3}(t,x)$ $=0\qquad(t,x)\in\Omega=[0,1]\times[0,1]$
$u(t,0)$ $=0$
$u(t,1)$ $=\cos(5\pi t)+t^{3}$
$u(0,x)$ $=x$
$u_{t}(0,x)$ $=0$
To solve this nonlinear PDE using the proposed QCPINN, we construct a composite loss function incorporating the physical constraints, boundary conditions, and initial conditions:
[…]
The loss weights are chosen empirically to balance the different components as $\lambda_{1}=1.0$, $\lambda_{2}=10.0$, and $\lambda_{3}=1.0$.
A.5 Convection-diffusion Equation
The problem focuses on solving the 2D convection-diffusion equation, a partial differential equation that models the transport of a quantity under convection (bulk movement) and diffusion (spreading due to gradients). The specific form considered here includes a viscous 2D convection-diffusion equation of the form:
$\begin{split}u_{t}+c_{1}u_{x}+c_{2}u_{y}-D\Delta u(x,y)&=0\qquad(t,x,y)\in\Omega\\ u(t,\mathbf{x})&=g_{0}(t,\mathbf{x})\qquad(t,\mathbf{x})\in\Gamma_{0}\\ u(t,\mathbf{x})&=g_{1}(t,\mathbf{x})\qquad(t,\mathbf{x})\in\Gamma_{1}\\ u(0,\mathbf{x})&=h(0,\mathbf{x})\qquad\mathbf{x}\in\partial\Omega\times[0,T]\end{split}$ (8)
where $c_{1}$ is the convection velocity in the $x$ direction, $c_{2}$ is the convection velocity in the $y$ direction, $D$ is the diffusion coefficient, and $\Delta=\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}$ is the Laplacian operator. For our numerical investigation, we choose an exact solution describing a Gaussian pulse:
$u(t,x,y)=\exp(-100((x-0.5)^{2}+(y-0.5)^{2}))\exp(-t)$
The corresponding initial condition is:
$h(0,x,y)=\exp(-100((x-0.5)^{2}+(y-0.5)^{2}))$
We choose $c_{1}=1.0$, $c_{2}=1.0$, and $D=0.01$. The complete initial-boundary value problem on the unit cube domain becomes:
$\begin{split}u_{t}(t,x,y)+u_{x}(t,x,y)+u_{y}(t,x,y)-0.01\left(u_{xx}(t,x,y)+u_{yy}(t,x,y)\right)&=0\\ u(t,x,y)&=g(t,x,y)\\ u(0,x,y)&=h(x,y)\end{split}$
where $(t,x,y)\in\Omega=[0,1]\times[0,1]\times[0,1]$ To solve this problem using the proposed QCPINN model, we formulate a loss function that incorporates the physical governing equation, boundary conditions, and initial conditions:
$\begin{split}\mathcal{L}(\theta)&=\underset{\theta}{\text{min}}\big{(}\lambda_{1}\|\mathcal{L}_{\text{phy}}(\theta)\|_{\Omega}+\lambda_{2}\|\mathcal{L}_{\text{bc}}(\theta)\|_{\Gamma_{1}}+\lambda_{3}\|\mathcal{L}_{\text{ic}}(\theta)\|_{\Gamma_{0}}\big{)}\\ &=\underset{\theta}{\text{min}}\big{[}\lambda_{1}\text{min}(\|\underbrace{u_{\theta_{t}}(t,x,y)+u_{\theta_{x}}(t,x,y)+u_{\theta_{y}}(t,x,y)-0.01(u_{\theta_{xx}}(t,x,y)+u_{\theta_{yy}}(t,x,y))}_{\text{PDE residual}}\|_{\Omega})\\ &+\lambda_{2}\text{min}(\|\underbrace{u(t,x,y)-g_{1}(t,x,y)}_{\text{Boundary conditions}}\|_{\Gamma_{1}})+\lambda_{3}\text{min}(\|\underbrace{u_{\theta}(0,x,y)-h(x,y)}_{\text{Initial conditions}}\|_{\Omega_{0}})\big{]}\end{split}$
[…]